OpenStudy (anonymous):

How would I set this problem up? To find a distance across a lake, a surveyor locates points A, B, and C as shown. What is the measure of angle B to the nearest degree? (Enter only the number.) a=108 b= 55 C=59 degrees. I know I can do the math, I just can’t remember how to set it up.

OpenStudy (ranga):

You have to show us the diagram. Use the "Draw" button below the reply box.

OpenStudy (anonymous):



OpenStudy (ranga):

Use the law of cosines to find AB Use the law of sines to find angle B

OpenStudy (ranga):

Law of cosines: c^2 = a^2 + b^2 - 2ab(cos(C))

OpenStudy (anonymous):

so side AB would be 92.57?

OpenStudy (ranga):

Law of sines: sin(B) / b = sin(C) / c

OpenStudy (anonymous):

and angle B would be 30.6 but rounded would be 31?

OpenStudy (ranga):

side AB = 92.58 is correct. I don't have a scientific calculator so I will have to do it online. Let me find the angle.

OpenStudy (ranga):

Yes angle B = 30.6 and round it up to 31. Your calculations are all correct.

OpenStudy (anonymous):

thank you much appreciated!

OpenStudy (ranga):

you are welcome.